# PredIN: Towards Open-Set Gesture Recognition via Prediction Inconsistency

### v1

@@ 5.1 Image Domain Verification
Fig. 7: The visualization of class feature distribution of two ensemble members. Both horizontal and vertical coordinates represent class labels. Each value represents the degree of proximity between two classes within a single member. (a) represents the two members of the Deep Ensemble model. There are similar local structures between the two members (yellow boxes). (b) represents the two members of the PredIN model. The global layout and local neighboring pairs are different (red boxes).
5.2 Ensemble Diversity Analysis
@@
The improvement in performance compared to the Deep Ensemble model confirms the increased ensemble diversity introduced by PredIN. To further evaluate the diversity, we visualize the class feature distribution and measure the diversity with a diversity metric $div$.
@@
We first verify that our approach achieves class feature distribution inconsistency. Specifically, we visualize the class feature distribution using a proximity matrix. As the class feature distribution can be approximately characterized by learned prototypes, we compute the distances between class prototypes and convert them into probabilities in order to represent class proximity. Fig. 7 demonstrates that PredIN achieves our desired class feature distribution inconsistency between ensemble members, especially compared to the Deep Ensemble model. Locally, each class has different neighboring classes, while globally, the relative positions of classes vary, which causes the feature spaces of the two members different.
Fig. 8: Diversity comparison on four sEMG datasets.
@@
To further validate that maximizing the class feature distribution inconsistency enhances ensemble diversity, we compare the PredIN to the Deep Ensemble model over a diversity metric. Common diversity metrics [14, 15] employed in classification tasks focus on the fraction of label changes in known samples and are therefore not directly applicable to OSR. Consequently, we modify these metrics to measure the relative prediction inconsistency between ensemble members on unknown samples:
$div=\frac{Fraction\ of\ unknown\ label\ changes}{Fraction\ of\ known\ label\ changes}.$ (18)
@@
The results in Fig. 8 demonstrate the consistent diversity improvements achieved by our approach across four sEMG datasets.
5.3 Ensemble Number
@@
In our experiments, we use two ensemble members because the inconsistency loss has a symmetrical form, which makes it unsuitable for optimizing more members in parallel within the ensemble. To evaluate whether more ensemble members will perform better, we train multiple models sequentially. The training of the first model is only based on the PL loss $\mathcal{L}_{PL}$, minimizing its empirical classification risk. Subsequent models apply our proposed approach and are optimized based on the loss $\mathcal{L}_{Div}$. Each model obtains a different class feature distribution from the former. Fig. LABEL:fig10 shows the results for a larger ensemble number of $5$ on bioDB2 and DB7. The more ensemble models gain better performance, which aligns with the assumption of ensemble learning. However, ensembling more members will lead to performance saturation and increased computational burden.
5.4 Limitations and Future Work
@@
Although our proposed PredIN addresses the challenge of open-set gesture recognition and demonstrates promising performance, it has several limitations. One limitation is the computation cost associated with the ensemble learning framework. Ensemble learning brings diversity but also introduces certain computational costs. In future work, we can mitigate this by sharing a shallow encoder among ensemble members. Additionally, enhancing the diversity of the ensemble and improving the performance of individual members are orthogonal objectives. By developing an individual member with optimal performance and combining it with the diversity of the ensemble model, we can further enhance the model’s rejection performance. The potential of ensemble models in OSR tasks remains to be fully explored. Moreover, our approach is limited to the rejection of unknown gestures, without further considering the system’s ability to learn and classify novel gestures dynamically. One potential future work is to incorporate open-world recognition frameworks, such as leveraging incremental learning techniques.
6 Conclusion
@@
Generalizing gesture recognition from closed-set to open-set is important for real-world HMI. To tackle open-set gesture recognition based on sEMG, we propose an ensemble learning approach, PredIN, based on our observed prediction inconsistency for the unknown due to the ensemble diversity. Specifically, we propose two complementary losses to improve OSR performance by enhancing ensemble diversity while maintaining individual performance. Extensive experiments conducted on multiple datasets consistently demonstrate that our approach outperforms previous state-of-the-art open-set classifiers. This means that our gesture recognition system can maintain high classification accuracy for predefined gestures, while effectively rejecting gestures of disinterest. We hope this work will boost the applications of gesture recognition technologies in real-world scenarios. Moving forward, we will also explore extending our technology to adapt to diverse recognition tasks.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to improve language and readability. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Acknowledgments
This work was partially supported by OYMotion Technologies.

### v2

@@ IV-D Comparison with the State-of-the-arts
Figure 8: The visualization of class feature distribution of dual perspectives. Both horizontal and vertical coordinates represent class labels. Each value represents the degree of proximity between two classes within a single branch. (a) represents the two branches of the Dual-perspective model. There are similar local structures between the two branches (yellow boxes). (b) represents the two branches of the PredIN model. The global layout and local neighboring pairs are different (red boxes).
IV-E Ablation Study
Module Ablation. As presented in Table III, each component within our method has been systematically integrated into the PL baseline to verify its necessity. We first introduce the Dual-perspective framework, which combines two identical networks with different random initializations. As shown in the second raw, the introduction of dual perspectives yields notable improvements in AUC and OSCR scores compared to the PL baseline, demonstrating the effectiveness of using prediction inconsistency for unknown rejection. Secondly, we further verify the effectiveness of our proposed two losses. In PredIN, the roles of inconsistency loss and triplet loss are complementary: the loss $\mathcal{L}_{trip}$ acts to optimize the inter-class separability within each branch while the inconsistency loss $\mathcal{L}_{incon}$ maximizes the class feature distribution inconsistency among different perspectives. To further explain, we apply the loss $\mathcal{L}_{trip}$ to the Dual-perspective model alone and then combine these two losses into the model training. The standalone application of $\mathcal{L}_{trip}$ enhances the individual rejection capability, thereby consistently improving overall performance compared to the Dual-perspective model. When combining these two losses, PredIN magnifies the prediction inconsistency by enhancing the differences between dual perspectives while maintaining the individual performance, which provides further improvements and clearly demonstrates the effectiveness of maximizing the class feature distribution inconsistency. In addition, we remove the loss $\mathcal{L}_{trip}$ from the PredIN to verify the importance of maintaining individual performance from another aspect. A clear decrease occurs on BioPat DB2 and Ninapro DB2 after the removal. Finally, incorporating all the above components improves AUC by an average of +3.8% compared to the baseline PL model, which demonstrates that each component contributes to the overall improvement on unknown rejection.
Hyperparameters Ablation. We evaluate the effect of two sets of hyperparameters: the trade-off weights $\beta$ and $\gamma$ in (6) and (12), and two margins $m_{1}$ and $m_{2}$ in (10) and (11). These experiments are conducted using the bioDB2 dataset. Figure 7 shows the AUC for different $\beta$ and $\gamma$ values. The weight $\beta$ influences the compactness of feature space in prototype learning. According to the results, larger values yield better results, but excessively large values may cause optimization issues and lead to a performance decrease. The weight $\gamma$ represents the degree of adjustment to the class feature distribution. Larger values similarly affect the model’s classification. Setting $\beta$ and $\gamma$ to $1.0$ is optimal. Two margins influence the degree of class feature distribution inconsistency and inter-class separability. Greater values of $m_{1}$ and $m_{2}$ lead to reduced inconsistency but increased enhancement for inter-class separability. A balanced result is that the optimal values for $m_{1}$ and $m_{2}$ are set to $0.5$ and $1.0$ in our experiments, respectively.
V Further Analysis and Discussion
V-A Dual-perspective Prediction Inconsistency Analysis
@@
The improvement in performance compared to the Dual-perspective model confirms the increased prediction inconsistency introduced by PredIN. To further evaluate this improvement, we visualize the class feature distribution and measure the prediction inconsistency with an inconsistency metric $Incon$.
@@
We first verify that our approach enlarges the differences between perspectives by maximizing the class feature distribution inconsistency. Specifically, we visualize the class feature distribution using a proximity matrix. As the class feature distribution can be approximately characterized by learned prototypes, we compute the distances between class prototypes and convert them into probabilities in order to represent class proximity. Figure 8 demonstrates that PredIN achieves our desired class feature distribution inconsistency between two branches, especially compared to the Dual-perspective model. Locally, each class has different neighboring classes, while globally, the relative positions of classes vary, which causes the feature spaces of the two perspectives different.
Figure 9: Inconsistency comparison on four sEMG datasets.
@@
To further validate that maximizing class feature distribution inconsistency enhances prediction inconsistency for unknown samples between two perspectives, we compare the PredIN with the Dual-perspective model using an inconsistency metric. We focus on the fraction of prediction changes both in known and unknown samples to measure the relative prediction inconsistency between perspectives on unknown samples:
$Incon=\frac{Fraction\ of\ unknown\ prediction\ changes}{Fraction\ of\ known\ prediction\ changes}.$ (17)
@@
The results in Figure 9 demonstrate the consistent prediction inconsistency improvements achieved by PredIN across four sEMG datasets.
V-B Perspective Number
@@
In our experiments, we use two perspectives because the inconsistency loss has a symmetrical form, which makes it unsuitable for optimizing more branches in parallel. To evaluate whether more perspectives will perform better, we train multiple models sequentially. The training of the first model is only based on the PL loss $\mathcal{L}_{PL}$, minimizing its empirical classification risk. Subsequent models apply our proposed approach and are optimized based on the loss $\mathcal{L}_{Div}$. Each model obtains a different class feature distribution from the former. Figure 10 shows the results for a larger perspective number of $5$ on BioPat DB2 and Ninapro DB7. The more perspectives gain better performance, which aligns with the prediction inconsistency among different perspectives. However, combining more perspectives will lead to performance saturation and increased computational burden.
Figure 10: Evaluation on the impact of number of perspectives. (a) BioPat DB2. (b) Ninapro DB7.
V-C Limitations and Future Work
@@
Although our proposed PredIN addresses the challenge of open-set gesture recognition and demonstrates promising performance, it has several limitations. One limitation is the computation cost associated with the dual-perspective framework. Dual-perspective inconsistency learning brings prediction inconsistency but also introduces certain computational costs. In future work, we can mitigate this by sharing a shallow encoder among branches and can combine multiple perspectives. Additionally, enhancing the differences among multiple perspectives and improving the performance of individual branches are orthogonal objectives. By developing an individual branch with optimal performance and combining it with the differences of the multi-perspective model, we can further enhance the model’s OSR performance. The potential of multi-perspective models in OSR tasks remains to be fully explored. Moreover, our approach is limited to the rejection of unknown gestures, without further considering the system’s ability to learn and classify novel gestures dynamically. One potential future work is to incorporate open-world recognition frameworks, such as leveraging incremental learning techniques.
VI CONCLUSION
@@
Generalizing gesture recognition from closed-set to open-set is important for real-world HMI. To tackle open-set gesture recognition based on sEMG, we propose a novel dual-perspective inconsistency learning approach, PredIN, based on our observed pronounced prediction inconsistency within different perspectives for the unknown. Specifically, we propose two complementary strategies to improve OSR performance by enhancing perspective differences while maintaining individual performance. Extensive experiments conducted on multiple public datasets consistently demonstrate that our approach outperforms previous state-of-the-art open-set classifiers. Compared to existing methods for closed-set gesture recognition systems, our approach can maintain high classification accuracy for predefined gestures, while effectively rejecting gestures of disinterest. We hope this work will boost the applications of gesture recognition technologies in real-world scenarios. Moreover, this research offers new insights into open-set recognition and has great potential to be extended to various fields. Moving forward, we will also explore extending our technology to adapt to diverse pattern recognition tasks.